# Measuring Diversity in Synthetic Datasets

### v2

@@ 1 Introduction
Since the diversity evaluation of synthetic datasets remains under-explored, a natural solution is to directly employ diversity evaluation methods from related fields, such as natural language processing (NLP) (Khurana et al., 2023) and machine learning (ML) (Jordan & Mitchell, 2015). Specifically, efforts to measure diversity within these domains can be summarized into three categories: N-gram-based method (Zhu et al., 2018; Mishra et al., 2020), Reference-based method (Heusel et al., 2017; Cífka et al., 2018), and Transformation-based method (Du & Black, 2019; Zhang et al., 2024). The n-gram-based method evaluates diversity through n-gram statistics, e.g., the distinct-n metric (Li et al., 2015) calculates the proportion of unique n-grams out of the total number of n-grams. This approach primarily focuses on the form differences of evaluated texts, often overlooking semantic aspects. To align the diversity criteria with human judgment, the reference-based method has emerged as a promising alternative. This approach employs a reference distribution or data as an approximation of human judgment and calculates the similarity between the evaluated data and the reference data (Holtzman et al., 2019). However, collecting reference data can be both time-consuming and may introduce potential biases.
@@ 3.1 LLM as a Dataset Generator
It is worth noting that not all components in $\mathcal{D}$ are generated by $\mathcal{M}$, which is related to the category of $\mathcal{D}_{sup}$. As shown in Table 1, $\mathcal{D}_{sup}$ can be divided into three categories, namely input text, output text, and seed samples. In Table 1, $\rightarrow$” and $\mathcal{T}_{seed}$ represent the direction of generation and seed samples, respectively. For example, “$x_{i}\rightarrow y_{i}$” signifies that, given input text $x_{i}$ denoted as $\mathcal{D}_{sup}$, $\mathcal{M}$ processes $\mathcal{D}_{sup}$ and $T$, generating the output text $y_{i}$.
@@ 4.1 DCScore: Measuring Diversity from a Classification Perspective
$=\Phi(\{\mathcal{\tilde{T}}_{i}\}_{i=1}^{n}),\\ \textbf{K}=\operatorname{Kernel}(\textbf{H}),$ H ) , (3)
@@ 5.2.3 Correlation with LLM Evaluator
To further verify the effectiveness of DCScore, we investigate the evaluation correlation between DCScore and LLMs. Following the setting in Section 5.2.2, we employ GPT-4 to conduct pairwise comparisons between two generated datasets with different $\tau_{g}$. These generated datasets are identical to those used in Section 5.2.2. Based on the pairwise comparison results, we obtain the diversity ranking outcomes. Regarding GPT-4 evaluation results as the diversity pseudo-truth, we report the pairwise evaluation correlation between DCScore, GPT-4, and $\tau_{g}$ in Table 5. We observe that DCScore exhibits strong correlations with GPT-4 and $\tau_{g}$ in zero-shot settings. By comparing the results of “$\tau_{g}$-DCScore” and “$\tau_{g}$-GPT-4”, we find that DCScore outperforms the GPT-4 evaluator in terms of correlation with $\tau_{g}$ in zero-shot settings. Regarding the correlation performance in few-shot settings, we notice lower correlations of all baseline methods compared to zero-shot settings. We guess that this phenomenon is related to the distributions of the generated datasets. Although DCScore exhibits lower correlations (about 0.6) with GPT-4, this result can still be considered a strong correlation according to (Akoglu, 2018).
In summary, DCScore exhibits a strong correlation (Spearman’s $\rho$geq 0.6), with three diversity pseudo-truths: $\tau_{g}$, human judgment, and LLMs evaluation, thereby verifying the effectiveness of DCScore.
@@ C.2 Implementation Details
For three transformation-based methods, including DCScore, VendiScore, and K-means Inertia, we employ unsup-simcse-bert-base-uncased (princeton nlp, 2021) as the weight of the embedding function. For all language models used to generate the dataset, we set the top-p and top-k parameters to 1 and -1, respectively. Additionally, we limit the maximum number of newly generated tokens to 100 for the text classification task and 30 for the story completion task. All experiments are conducted on 8$\times$ NVIDIA Tesla V100 GPU with 32GB of memory. For self-generated datasets, we only evaluate the diversity of generated components.
@@ Appendix E Baseline Methods
[…]

### v3

@@ 1 Introduction
Since the diversity evaluation of synthetic datasets remains under-explored, a natural solution is to directly employ diversity evaluation methods from related fields, such as natural language processing (NLP) (Khurana et al., 2023) and machine learning (ML) (Jordan & Mitchell, 2015). Specifically, efforts to measure diversity within these domains can be summarized into three categories: N-gram-based method (Zhu et al., 2018; Mishra et al., 2020), Reference-based method (Heusel et al., 2017; C´ıfka et al., 2018), and Transformation-based method (Du & Black, 2019; Zhang et al., 2024). The n-gram-based method evaluates diversity through n-gram statistics, e.g., distinct-n (Li et al., 2015), focusing on textual form rather than semantic content. To align the diversity criteria with human judgment, the reference-based method has emerged as a promising alternative. This approach employs a reference distribution or data as an approximation of human judgment and calculates the similarity between the evaluated data and the reference data (Holtzman et al., 2019). However, collecting reference data can be both time-consuming and may introduce potential biases.
@@ 3.1 LLM as a Dataset Generator
It is worth noting that not all components in $\mathcal{D}$ are generated by $\mathcal{M}$, which is related to the category of $\mathcal{D}_{sup}$. As shown in Table 1, $\mathcal{D}_{sup}$ can be divided into three categories, namely input text, output text, and seed samples. In Table 1, $\rightarrow$” and $\mathcal{T}_{seed}$ represent the direction of generation and seed samples, respectively. For example, $x_{i}\rightarrow y_{i}$” signifies that, given input text $x_{i}$ denoted as $\mathcal{D}_{sup}$, $\mathcal{M}$ processes $\mathcal{D}_{sup}$ and $T$, generating the output text $y_{i}$.
@@ 4.1 DCScore: Measuring Diversity from a Classification Perspective
$\textbf{H}=\Phi(\{\mathcal{\tilde{T}}_{i}\}_{i=1}^{n}),\\ \textbf{K}=\operatorname{Kernel}(\textbf{H}),$ (3)
@@ 5.2.3 Correlation with LLM Evaluator
To further verify the effectiveness of DCScore, we investigate the evaluation correlation between DCScore and LLMs. Following the setting in Section 5.2.2, we employ GPT-4 to conduct pairwise comparisons between two generated datasets with different $\tau_{g}$. These generated datasets are identical to those used in Section 5.2.2. Based on the pairwise comparison results, we obtain the diversity ranking outcomes. Regarding GPT-4 evaluation results as the diversity pseudo-truth, we report the pairwise evaluation correlation between DCScore, GPT-4, and $\tau_{g}$ in Table 5. We observe that DCScore exhibits strong correlations with GPT-4 and $\tau_{g}$ in zero-shot settings. By comparing the results of $\tau_{g}$-DCScore” and $\tau_{g}$-GPT-4”, we find that DCScore outperforms the GPT-4 evaluator in terms of correlation with $\tau_{g}$ in zero-shot settings. Regarding the correlation performance in few-shot settings, we notice lower correlations of all baseline methods compared to zero-shot settings. We guess that this phenomenon is related to the distributions of the generated datasets. Although DCScore exhibits lower correlations (about 0.6) with GPT-4, this result can still be considered a strong correlation according to (Akoglu, 2018).
In summary, DCScore exhibits a strong correlation (Spearman’s $\rho$ $\geq$ 0.6), with three diversity pseudo-truths: $\tau_{g}$, human judgment, and LLMs evaluation, thereby verifying the effectiveness of DCScore.
@@ C.2 Implementation Details
[…]
@@ Appendix E Baseline Methods
$\begin{split}&\textnormal{{Text Representation: }}\operatorname{n-grams}(\operatorname{Concat}(\mathcal{D})),\\ &\textnormal{{Pairwise Similarity: }}\operatorname{Unique}(\operatorname{n-grams}(\operatorname{Concat}(\mathcal{D}))),\\ &\textnormal{{Diversity Summarization: }}\operatorname{Distinct-n}(\mathcal{D})=\frac{|\operatorname{Unique}(\operatorname{n-grams}(\operatorname{Concat}(\mathcal{D})))|}{|\operatorname{n-grams}(\operatorname{Concat}(\mathcal{D}))|},\end{split}$ (18)